\begin{thebibliography}{99}


\bibitem{AS} G. Alessandrini and  J. Sylvester,
\newblock Stability for multidimensional inverse spectral problem, 
\newblock Commun. Partial Diffent. Equat. 15 (5) (1990), 711-736.

\bibitem{Bel87} M. Belishev,
\newblock An approach to multidimensional inverse problems for the wave equation, 
\newblock Dokl. Akad. Nauk SSSR 297 (1987), 524-527.

\bibitem{Bel92} M. Belishev and Y. Kurylev,
\newblock To the reconstruction of a Riemannian manifold via its spectral data (BC-method), 
\newblock Commun. Partial Diffent. Equat. 17 (1992), 767-804.

\bibitem{BCDKS} M. Bellassoued, M. Choulli, D. Dos Santos Ferreira, Y. Kian and P. Stefanov, 
\newblock A Borg-Levinson theorem for magnetic Schr\"odinger operators on a Riemannian manifold,
\newblock to appear in  Ann. Inst. Fourier.

\bibitem{BCKS} M. Bellassoued, M. Choulli,  Y. Kian and E. Soccorsi,
\newblock Borg Levinson type theorem with partial spectral boundary data,
\newblock preprint.

\bibitem{BCY} M. Bellassoued, M. Choulli and M. Yamamoto, 
\newblock Stability estimate for an inverse wave equation and a multidimensional Borg-Levinson theorem, 
\newblock J. Diffent. Equat. 247 (2) (2009), 465-494.

\bibitem{BCY2} M. Bellassoued, M. Choulli and M. Yamamoto, 
\newblock Stability estimate for a multidimensional inverse spectral problem with partial data, 
\newblock J. Math. Anal. Appl. 378 (1) (2011), 184-197.


\bibitem{CS} P. Caro and M. Salo, 
\newblock Stability of the Calder\'on problem in admissible geometries,
\newblock Inverse Problems and Imaging 8 (2014),  939-957.

\bibitem{Ch} M. Choulli,
\newblock Une introduction aux probl\`emes inverses elliptiques et paraboliques, 
\newblock Math\'ematiques et Applications, Vol. 65, Springer-Verlag, Berlin, 2009.

\bibitem{ChS} M. Choulli and P. Stefanov,
\newblock Stability for the multi-dimensional Borg-Levinson theorem with partial spectral data, 
\newblock Commun. Partial Diffent. Equat. 38 (3) (2013), 455-476.

\bibitem{IY} Oleg Yu. Imanuvilov and M. Yamamoto,
\newblock Stability of determination of Riemannian metrics by spectral data and Dirichlet-to-Neumann map limited on arbitrary subboundary,
\newblock Inverse Problems and Imaging 13 (6) (2019), 1213-1258.

\bibitem{Is} H. Isozaki,
\newblock Some remarks on the multi-dimensional Borg-Levinson theorem,
\newblock J. Math. Kyoto Univ. 31 (3) (1991), 743-753.

\bibitem{KK} A. Katchalov and Y. Kurylev,
\newblock Multidimensional inverse problem with incomplete boundary spectral data,
\newblock Commun. Partial Diffent. Equat. 23 (1998),  55-95.

\bibitem{KKL} A. Katchalov,  Y. Kurylev, M.  Lassas,
\newblock Inverse boundary spectral problems, 
\newblock Chapman \& Hall/CRC, Boca Raton, FL, 2001, 123, xx+290.

\bibitem{KKS} O. Kavian, Y. Kian, E. Soccorsi,
\newblock Uniqueness and stability results for an inverse spectral problem in a periodic waveguide, 
J. Math. Pures Appl. 104 (2015), no. 6, 1160-1189.

\bibitem{Ki1} Y. Kian,
\newblock A multidimensional Borg-Levinson theorem for magnetic Schr\"odinger operators with partial spectral data, 
\newblock J. Spectral Theory 8 (2018), 235-269.

\bibitem{KOM} Y. Kian, L. Oksanen, M. Morancey, 
\newblock Application of the boundary control method to partial data Borg-Levinson inverse spectral problem,
\newblock Math. Control and Related Fields 9 (2019),  289-312.

\bibitem{NSU} A. Nachman, J. Sylvester, G. Uhlmann,
\newblock An n-dimensional Borg-Levinson theorem, 
\newblock Commun. Math. Phys. 115 (4) (1988), 595-605.

\bibitem{Po} V. Pohjola, 
\newblock Multidimensional Borg-Levinson theorems for unbounded potentials
\newblock  Asymptot. Anal. 110 (3-4) (2018),  203-226.


\end{thebibliography}

\documentclass[english]{amsart}

\usepackage{esint}
\usepackage[svgnames]{xcolor} 
\usepackage[colorlinks,citecolor=red,pagebackref,hypertexnames=false,breaklinks]{hyperref}
\usepackage{pgf,tikz}
\usepackage{pdfsync}

\usepackage{dsfont}
\usepackage{url}
\usepackage[utf8]{inputenc}
\usepackage[T1]{fontenc}
\usepackage{lmodern}
\usepackage{babel}
\usepackage{mathtools}  
\usepackage{amssymb}
\usepackage{lipsum}
\usepackage{mathrsfs}
\usepackage{color}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{lemma}{Lemma}[section]

\newtheorem{corollary}{Corollary}[section]

\numberwithin{equation}{section}

\title[Borg Levinson type theorem]{A simple proof of a multidimensional Borg-Levinson type theorem}

\author[Choulli]{Mourad Choulli}
\address{Universit\'e de Lorraine, 34 cours L\'eopold, 54052 Nancy cedex, France}
\thanks{The author supported by the grant ANR-17-CE40-0029 of the French National Research Agency ANR (project MultiOnde).}
\email{mourad.choulli@univ-lorraine.fr}

\date{\today}

\subjclass[2010]{35R30, 35J10.} 
\keywords{Admissible Riemannian manifold, simple Riemannian manifold, boundary spectral data, Borg-Levinson type theorem, Dirichlet-to-Neumann map}

\begin{document}

\begin{abstract}
We provide a simple and short proof of a multidimensional Borg-Levinson type theorem. Precisely, we prove that the knowledge of spectral boundary data determine uniquely the corresponding potential  appearing in the Sch\"odinger operator on an admissible Riemannian manifold. We also sketch the proof of uniqueness in the case of incomplete spectral boundary data. The new results in the present work complete those existing in the literature.
\end{abstract}

\maketitle

\section{Introduction}\label{section1}

Let $\mathcal{M}=(\mathcal{M},\mathfrak{g})$ be a smooth compact Riemannian manifold, of dimension $n\ge3$,  with boundary $\partial \mathcal{M}$. We recall that, in local coordinates,  the usual Laplace-Beltrami operator is given by
\[
\Delta_\mathfrak{g}=\frac{1}{\sqrt{|\mathfrak{g}|}}\sum_{j,k=1}^n\frac{\partial}{\partial x_j}\left(\sqrt{|\mathfrak{g}|}\mathfrak{g}^{jk}\frac{\partial}{\partial x_k}\, \cdot \right).
\]

We define, for $q\in L^\infty (\mathcal{M},\mathbb{R})$,  the unbounded operator $L=L(q)$ acting on $L^2(\mathcal{M})$ as follows
\[
L=-\Delta _{\mathfrak{g}}+q\quad \mbox{with}\quad D\left(L \right)=H_0^1(\mathcal{M})\cap H^2(\mathcal{M}).
\]
As $L$ is self-adjoint operator with compact resolvent, its spectrum is reduced to a sequence of eigenvalues:
\[
-\infty < \lambda_1\le \ldots \lambda_k\le \ldots \quad \mbox{and}\quad \lambda_k\rightarrow \infty \; \mbox{as}\; k\rightarrow \infty.
\]
Of course $\lambda_k=\lambda_k (q)$, $k\ge 1$.

Furthermore, there exists $(\phi_k)$, $\phi_k=\phi_k(q)$, $k\ge 1$, an orthonormal basis of $L^2(\mathcal{M})$ consisting of eigenfunctions. Each $\phi_k$ is associated to $\lambda_k$, $k\ge 1$.

For simplicity convenience, we use the notation 
\[
\psi_k=\partial_\nu  \phi_k|{_{\partial \mathcal{M}}}, \quad k\ge 1,
\]
where $\nu$ is the unit normal exterior vector field on $\partial \mathcal{M}$ with respect to $\mathfrak{g}$.

When $q\in L_+^\infty (\mathcal{M})=\{q\in L^\infty (\mathcal{M},\mathbb{R});\; q\ge 0\}$, $0$ is not in the spectrum of $L$ and all the eigenvalues of $L$ are positive. Whence, for any $f\in H^{1/2}(\partial \mathcal{M})$, the BVP
\begin{equation}\label{bvp}
(-\Delta _{\mathfrak{g}}+q)u=0\; \mbox{in}\; \mathcal{M}\quad \mbox{and}\quad \mbox u=f\; \mbox{on}\; \partial \mathcal{M}
\end{equation}
admits a unique solution $u=u(q,f)\in H^1(\mathcal{M})$ with $\partial_\nu u\in H^{-1/2}(\partial \mathcal{M})$. Moreover 
\[
\|\partial_\nu u\|_{H^{-1/2}(\partial \mathcal{M})}\le C\|f\|_{H^{1/2}(\partial \mathcal{M})},
\]
where the constant $C$ is independent of $f$. In other words, the mapping
\[
\Lambda(q) :f\in H^{1/2}(\partial \mathcal{M})\rightarrow \partial_\nu u\in H^{-1/2}(\partial \mathcal{M})
\]
defines a bounded operator. The operator $\Lambda (q)$ is usually called the Dirichlet-to-Neumann map associated to $q$.

Pick $\tilde{q}\in L_+^\infty (\mathcal{M})$. Set $\tilde{\lambda}_k=\lambda_k (\tilde{q})$, $\tilde{\phi}_k=\phi_k (\tilde{q})$ and $\tilde{\psi}_k=\psi_k (\tilde{q})$. Introduce then the notation 
\[
\mathfrak{D}(q,\tilde{q})=\sum_{k\ge 1}\left|\lambda_k-\tilde{\lambda}_k\right|+\sum_{k\ge 1} \|\tilde{\psi}_k-\psi_k\|_{L^2(\partial \mathcal{M})}.
\]

We aim to prove the following result:
 
 \begin{theorem}\label{theorem1} 
 Let $\aleph >0$. Suppose that $0\le q\le \aleph$, $0\le \tilde{q}\le \aleph$ and $\mathfrak{D}(q,\tilde{q})<\infty$. Then $\Lambda (q) -\Lambda(\tilde{q})$ extends to a bounded operator on $L^2(\partial \mathcal{M})$ and
\[
\|\Lambda (q) -\Lambda(\tilde{q})\|_{\mathscr{B}(L^2(\partial \mathcal{M}))}\le C\mathfrak{D}(q,\tilde{q}),
\]
where the constant $C$ only depends on $n$, $\mathcal{M}$ and $\aleph$.
 \end{theorem}

Let $(\mathscr{M},g)$ a compact Riemannian manifold with boundary $\partial \mathscr{M}$. We say that $\mathscr{M}$ is admissible if $\mathscr{M}\Subset \mathbb{R}\times \mathscr{M}_0$, for some (n-1)-dimensional simple manifold $(\mathscr{M}_0,g_0)$,  and if $g=c(\mathfrak{e}\oplus g_0)$, where $\mathfrak{e}$ is the Euclidean metric on $\mathbb{R}$ and $c$ is a smooth positive function on $\mathscr{M}$. A compact Riemannian manifold with boundary $(\mathscr{M}_0,g_0)$ is called simple  if, for any $x\in \mathscr{M}_0,$ the exponential map $\exp_x$ with its maximal domain of definition is a diffeomorphism onto $\mathscr{M}_0$, and if $\partial \mathscr{M}_0$ is strictly convex (which means that the second fundamental form of $\partial \mathscr{M}_0\hookrightarrow \mathscr{M}_0$ is positive definite).

Observing that 
\[
\|\Lambda (q) -\Lambda(\tilde{q})\|_{\mathscr{B}(H^{1/2}(\partial \mathcal{M}),H^{-1/2}(\partial \mathcal{M}))}\le \|\Lambda (q) -\Lambda(\tilde{q})\|_{\mathscr{B}(L^2(\partial \mathcal{M}))},
\]
and taking into account that $\mathfrak{D}(q,\tilde{q})=\mathfrak{D}(q-\lambda ,\tilde{q}-\lambda)$, for any $\lambda$ not belonging to the spectrum of $L(q)$ and $L(\tilde{q})$, we deduce as a straightforward consequence of Theorem \ref{theorem1} and \cite[Theorem 1]{CS} the following stability inequality: 

\begin{corollary}\label{corollary1}
Let $\aleph>0$, $t\in (0,1/2)$ and assume that $\mathcal{M}$ is admissible with $\mathcal{M}\Subset \mathbb{R}\times \mathcal{M}_0$. Then there exist two constants $C>0$ and $0<c<1$, only depending on $n$, $\mathcal{M}$ and $\aleph$, so that, for any $q,\tilde{q}\in L^\infty (\mathcal{M},\mathbb{R}) \cap H^t(\mathcal{M})$ satisfying 
\[ 
\|q\|_{L^\infty(\mathcal{M})\cap H^t(\mathcal{M})}\le \aleph ,\quad \|\tilde{q}\|_{L^\infty(\mathcal{M})\cap H^t(\mathcal{M})}\le \aleph\quad  \mbox{and}\quad  \mathfrak{D}(q,\tilde{q})\le c, 
\]
we have
\[
\| q-\tilde{q}\|_{L^2(\mathbb{R},H^{-3}(\mathcal{M}_0))}\le C\left| \ln \left[\mathfrak{D}(q,\tilde{q})+\left|\ln \left( \mathfrak{D}(q,\tilde{q})\right)\right|^{-1}\right]\right|^{-t/4}.
\]
\end{corollary}

This corollary says  in particular that in an admissible manifold the spectral boundary data $(\lambda_k(q),\psi_k(q))_{k\ge 1}$ determine uniquely $q\in L^\infty(\mathcal{M},\mathbb{R})\cap H^t(\mathcal{M})$:

\begin{theorem}\label{theorem2}
Let $t\in (0,1/2)$ and assume that $\mathcal{M}$ is admissible. If $q,\tilde{q}\in L^\infty(\mathcal{M},\mathbb{R})\cap H^t(\mathcal{M})$ satisfy
\[
\lambda_k(q)=\lambda_k(\tilde{q})\quad \mbox{and}\quad \psi_k(q)=\psi_k(\tilde{q}),\quad k\ge 1,
\]
then $q=\tilde{q}$.
\end{theorem}

Results of the same kind as in Theorem \ref{theorem2} are known in the literature as multidimensional Borg-Levinson type theorems.

The paper by Nachman, Sylvester and Uhlmann \cite{NSU} was the starting point of new developments and results on  multidimensional Borg-Levinson type theorems. We quote here the following short list of references: \cite{AS, Bel87, Bel92, BCDKS, BCY,BCY2,Ch,ChS,Is,KK,KKL,KKS,Ki1,KOM}. Of course many other works on spectral inverse problems exist in the literature. A  short survey on multidimensional Borg-Levinson type theorems can be found in \cite{BCDKS}. The case of unbounded potentials was considered by Pohjola in  \cite{Po}. We also mention a new stability inequality in two dimensional case, recently  established by Imanuvilov and Yamamoto in \cite{IY}.

\section{Proof of the main result}

Henceforward, the usual scalar products on $L^2(\mathcal{M})$ and $L^2(\partial \mathcal{M})$ are denoted respectively by $(\cdot |\cdot )$ and $\langle \cdot |\cdot \rangle$.

The following lemma will be useful in the sequel.
\begin{lemma}\label{lemma0}
Let $q\in L^\infty_+(\mathcal{M})$, $f\in H^{1/2}(\mathcal{M})$ and $u=u(q,f)$. We have
\begin{equation}\label{eq0}
u=-\sum_{k\ge 1}\frac{\langle f|\psi_k\rangle}{\lambda_k}\phi_k
\end{equation}
and hence
\[
\sum_{k\ge 1}\frac{|\langle f|\psi_k\rangle|^2}{\lambda_k^2}=\|u\|_{L^2(\mathcal{M})}^2.
\]
\end{lemma}

\begin{proof}
As $(\phi_k)$ is an orthonormal basis of $L^2(\mathcal{M})$, we obtain
\begin{equation}\label{eq2}
u=\sum_{k\ge 1}(u|\phi_k)\phi_k.
\end{equation}
Using that $u$ is the solution of the BVP \eqref{bvp} and Green's formula in order to get
\begin{align*}
(u|\lambda_k\phi_k)&=(u| (-\Delta_{\mathfrak{g}}+q)\phi_k)
\\
&= ((-\Delta_{\mathfrak{g}}+q)u|\phi_k)-\langle u|\partial_\nu \phi_k\rangle
\\
&=-\langle f|\partial_\nu \phi_k\rangle .
\end{align*}
That is we have
\[
(u|\phi_k)=-\frac{\langle f|\psi_k\rangle}{\lambda_k}.
\]
This identity in \eqref{eq2} yields \eqref{eq0}.
\end{proof}

Pick $(\rho_k)$ an orthonormal basis of $L^2(\mathcal{M} )$, $s\ge 0$ and define 
\[
\mathscr{H}^s=\left\{w\in H^2(\mathcal{M} ),\; \sum_{k\ge 1}k^{4s/n}\left(k^{4/n}|(w|\rho_k)|^2+|(\Delta_\mathfrak{g}w|\rho_k)|^2\right)<\infty\right\}.
\]
We have by Parseval's identity
\begin{align*}
&\sum_{k\ge 1}k^{4s/n}\left(k^{4/n}|(w|\rho_k)|^2+|(\Delta_\mathfrak{g}w|\rho_k)|^2\right)
\\
&\qquad =\left\| \sum_{k\ge 1} k^{2(s+1)/n}(w|\rho_k)\rho_k\right\|_{L^2(\mathcal{M})}^2
+\left\| \sum_{k\ge 1} k^{2s/n}(\Delta_{\mathfrak{g}}w|\rho_k)\rho_k\right\|_{L^2(\mathcal{M})}^2.
\end{align*}
In consequence, the left hand side of this identity does not depend on the choice of the orthonormal basis $(\rho_k)$. We endow $\mathscr{H}^s$ with the norm
\[
\|w\|_{\mathscr{H}^s}=\|w\|_{H^2(\mathcal{M})}+\left(\sum_{k\ge 1}k^{4s/n}\left(k^{4/n}|(w|\rho_k)|^2+|(\Delta_\mathfrak{g}w|\rho_k)|^2\right)\right)^{1/2}.
\]

It is not hard check that $\mathscr{H}^s$ contains all the eigenfunctions $(\varphi_k)$ of the operator $-\Delta_{\mathfrak{g}}$, under Neumann boundary condition, and therefore it contains also the closure, with respect to $\|\cdot \|_{\mathscr{H}^s}$, of the vector space spanned by these eigenfunctions. This can be easily seen just  by taking $\rho_k=\varphi_k$, $k\ge 1$, in the definition of $\mathscr{H}^s$.

We associate to $\mathscr{H}^s$ the  trace space
\[
\mathfrak{H}^s=\{ g\in H^{3/2}(\partial \mathcal{M});\; g=w|_{\partial \mathcal{M}}\;  \mbox{for some}\; w\in \mathscr{H}^s\}
\]
that we equip  with its natural quotient norm
\[
\|g\|_{\mathfrak{H}^s}=\inf\left\{\|w\|_{\mathscr{H}^s};\; w|_{\partial \mathcal{M}}=g\right\}.
\]

\begin{lemma}\label{lemma1}
We fix $s>n/4$. Then, for any $g\in \mathfrak{H}^s$, the series $\sum_{k\ge 1}\frac{\langle g|\psi_k\rangle}{\lambda_k}\phi_k $ converges in $H^2(\mathcal{M})$ with
\begin{equation}\label{1}
\sum_{k\ge 1}\frac{|\langle g|\psi_k\rangle|}{\lambda_k}\|\phi_k \|_{H^2(\mathcal{M})}\le C\|g\|_{\mathfrak{H}^s},
\end{equation}
where the constant $C$ is independent of $g$.
\end{lemma}

\begin{proof}
Let $g\in \mathfrak{H}^s$ and $w\in \mathscr{H}^s$ arbitrary so that $w|_{\partial \mathcal{M}}=g$. From Green's Formula, we have
\begin{align*}
\langle g|\psi_k\rangle =\langle g|\partial _\nu \phi_k\rangle &=(w|\Delta_\mathfrak{g}\phi_k)-(\Delta_\mathfrak{g}w|\phi_k).
\\
&= (w|(-\lambda_k+q)\phi_k )-(\Delta_\mathfrak{g}w|\phi_k ).
\end{align*}
Therefore, taking into account that $\lambda_k\sim k^{2/n}$ as $k\rightarrow +\infty$ (by Weyl's asymptotic formula), we get
\[
|\langle g|\psi_k\rangle|\le C\left( k^{2/n}|(\phi_k|w)|+|(\phi_k|\Delta_\mathfrak{g}w)| \right),\quad k\ge 1,
\]
where the constant $C$ is independent of $w$. 

But $\|\phi_k\|_{H^2(\mathcal{M})}\le C\lambda_k$ for any $k\ge 1$. Hence
\[
\frac{|\langle g|\psi_k\rangle|}{\lambda_k}\|\phi \|_{H^2(\mathcal{M})}\le Ck^{-2s/n}\left[k^{2s/n}\left( k^{2/n}|(\phi_k|w)|+|(\phi_k|\Delta_\mathfrak{g}w)| \right)\right],\quad k\ge 1.
\]
Whence, the series $\sum_{k\ge 1}\frac{\langle g|\psi_k\rangle}{\lambda_k}\phi_k $ converges in $H^2(\mathcal{M})$ and using Cauchy-Schwarz's inequality  we find
\[
\sum_{k\ge 1}\frac{|\langle g|\psi_k\rangle|}{\lambda_k}\|\phi_k \|_{H^2(\mathcal{M})}\le  C\|w\|_{\mathscr{H}^s}.
\]
As $w\in \mathscr{H}^s$ is chosen arbitrary so that $w|_{\partial \mathcal{M}}=g$, inequality \eqref{1} follows.
\end{proof}

\begin{proof}[Proof of Theorem \ref{theorem1}]
In this proof $C$ is a generic constant depending only on $n$, $\mathcal{M}$ and $\aleph$.

In light of Lemma \ref{lemma0} and Lemma \ref{lemma1}, we have
\[
\Lambda(q)(f)= -\sum_{k\ge 1}\frac{\langle f|\psi_k\rangle}{\lambda_k}\psi_k,\quad f\in \mathfrak{H}^s
\]
and
\[
\Lambda(\tilde{q})(f)= -\sum_{k\ge 1}\frac{\langle f|\tilde{\psi}_k\rangle}{\tilde{\lambda}_k}\tilde{\psi}_k,\quad f\in \mathfrak{H}^s.
\]

We split $\Lambda_q(f)-\Lambda_{\tilde{q}}(f)$, $f\in \mathfrak{H}^s$, into three terms:
\[
\Lambda_q(f)-\Lambda_{\tilde{q}}(f)=A_1+A_2+A_3,
\]
with
\begin{align*}
&A_1= \sum_{k\ge 1}\left(\frac{1}{\tilde{\lambda}_k}-\frac{1}{\lambda_k}\right)\langle f|\tilde{\psi}_k\rangle\tilde{\psi}_k,
\\
&A_2=\sum_{k\ge 1}\frac{\langle f|\tilde{\psi}_k\rangle-\langle f|\psi_k\rangle}{\lambda_k}\tilde{\psi}_k,
\\
&A_3=\sum_{k\ge 1}\frac{\langle f|\psi_k\rangle}{\lambda_k}\left(\tilde{\psi}_k-\psi_k\right).
\end{align*}

Then it is straightforward to check that
\begin{align*}
&\|A_1\|_{L^2(\partial \mathcal{M})}\le C\sum_{k\ge 1}\left|\lambda_k-\tilde{\lambda}_k\right|\|f\| _{L^2(\partial \mathcal{M})},
\\
&\|A_2\|_{L^2(\partial \mathcal{M})}+\|A_3\|_{L^2(\partial \mathcal{M})}\le C\sum_{k\ge 1} \|\tilde{\psi}_k-\psi_k\|_{L^2(\partial \mathcal{M})}\|f\| _{L^2(\partial \mathcal{M})}.
\end{align*}

Therefore, under the assumption $\mathfrak{D}(q,\tilde{q})<\infty$, the operator $\Lambda_q-\Lambda_{\tilde{q}}$ extends to a bounded operator on $L^2(\partial \mathcal{M})$  with
\[
\|\Lambda_q-\Lambda_{\tilde{q}}\|_{\mathscr{B}(L^2(\partial \mathcal{M}))}\le C\mathfrak{D}(q,\tilde{q}).
\]
The proof is then complete.
\end{proof}

\section{The case of incomplete spectral boundary data}

We explain briefly in this short section how we can get a Borg-Levinson type theorem with incomplete spectral boundary data. For this purpose, we assume that $\mathcal{M}$ is simple. In that case it is known that there exists another simple manifold $\mathcal{M}_1$ so that   $\mathcal{M}_1\Supset \mathcal{M}$. 

Fix $\aleph >0$. Let $q, \tilde{q}\in L^\infty (\mathcal{M},\mathbb{R})$ satisfying  $p=(q-\tilde{q})\chi_{\mathcal{M}}\in H^1(\mathcal{M}_1)$,
\[
\|q\|_{L^\infty(\mathcal{M})}\le \aleph ,\quad \|\tilde{q}\|_{L^\infty(\mathcal{M})}\le \aleph\quad  \mbox{and}\quad \|p\|_{H^1(\mathcal{M}_1)}\le \aleph. 
\]
From \cite{BCKS}, we have
\begin{equation}\label{2}
\|p\|_{L^2(\mathcal{M})}^4\le C\left(\frac{1}{|\Im \lambda |}+|\Im\lambda|\|\Lambda_{q-\lambda }- \Lambda_{\tilde{q}-\lambda } \|_{\mathscr{B}(L^2(\partial \mathcal{M} ))}\right),\quad \lambda \in \mathbb{C}\setminus \mathbb{R},
\end{equation}
where the constant $C$ only depends on $n$, $\mathcal{M}$ and $\aleph$.

Now if, for some fixed integer $\ell \ge 1$, $(\lambda_k, \psi_k)=(\tilde{\lambda}_k, \tilde{\psi}_k)$, $k\ge \ell$, then the calculations in the preceding section yield 
\begin{equation}\label{3}
\|\Lambda_{q-\lambda }- \Lambda_{\tilde{q}-\lambda } \|_{\mathscr{B}(L^2(\partial \mathcal{M} ))}\le C\left(\sum_{k=1}^\ell \frac{\lambda_k^2}{|\lambda _k-\lambda|}+\sum_{k=1}^\ell \frac{\tilde{\lambda}_k^2}{|\tilde{\lambda} _k-\lambda|} \right).
\end{equation}
Again, the constant $C$ only depends on $n$, $\mathcal{M}$ and $\aleph$.

We deduce by combining together \eqref{2} and \eqref{3} that there exist two constants $C>0$ and $\kappa >0$, only depending on  $n$, $\mathcal{M}$, $\aleph$ and $\ell$, so that 
\begin{equation}\label{4}
\|p\|_{L^2(\mathcal{M})}^4\le C\left(\frac{1}{|\Im \lambda |}+\frac{|\Im\lambda|}{|\lambda|}\right),\quad \lambda \in \mathbb{C}\setminus \mathbb{R},\; |\lambda|\ge \kappa.
\end{equation}

We find by taking in this inequality $\lambda =(\tau +i)^2=(\tau^2 -1)+2i\tau$, with $\tau \ge 1$ sufficiently large,
\[
\|p\|_{L^2(\mathcal{M})}^4\le C\left(\frac{1}{\tau }+\frac{\tau}{\tau ^2-1}\right).
\]
Taking the limit, as $\tau \rightarrow \infty$, in this inequality, we obtain $p=0$. In other words, we proved that the spectral boundary data $(\lambda_k,\psi_k)_{k\ge \ell}$ determine uniquely $q\in L^\infty (\mathcal{M},\mathbb{R})\cap H_0^1(\mathcal{M})$, for any arbitrary fixed integer $\ell \ge 1$:

\begin{theorem}\label{theorem3}
Let $\ell \ge 1$ be an integer and assume that the manifold $\mathcal{M}$ is simple. If $q,\tilde{q}\in L^\infty(\mathcal{M},\mathbb{R})\cap H_0^1(\mathcal{M})$ satisfy
\[
\lambda_k(q)=\lambda_k(\tilde{q})\quad \mbox{and}\quad \psi_k(q)=\psi_k(\tilde{q}),\quad k\ge \ell,
\]
then $q=\tilde{q}$.
\end{theorem}

\end{document}